### v4

# Origin of the hidden energy scale and the $f$-ratio in geometrically frustrated magnets

Abstract
Sufficiently clean geometrically frustrated (GF) magnets are the largest class of candidate materials that may host quantum spin liquids (QSLs). Some of them have been shown to exhibit spin-glass freezing, potentially precluding QSLs, at the “hidden energy scale”, which is significantly lower than the microscopic energy scale of spin interactions. Here, we investigate the origin of the hidden energy scale and its relationship to the $f$-ratio, the figure of merit for the degree of frustration in GF magnetic materials. The available experimental and numerical data provide evidence that GF magnets display, universally, two distinct temperature scales in the specific heat, the lowest of which is of the order of the hidden energy scale $T^{*}$. We argue that this scale is determined by non-magnetic excitations, similar to spin exchanges in chains of spins. The collective entropy of such excitations matches the entropy of the ground states of the Ising model on the same lattice, which provides a way to verify the proposed scenario in experiment. We demonstrate that in the presence of quenched disorder, a broad class of materials exhibits spin-glass freezing at temperatures of order $T^{*}$, in accordance with experimental observations. As $T^{*}$ is a property of the clean GF medium, it leads to a constraint on the $f$-ratio.
Geometrically frustrated (GF) magnets are considered to be a promising platform for hosting coveted quantum spin-liquid (QSL) states [43]. The geometry of the lattice combined with antiferromagnetic interactions leads to competing configurations of spins in such materials, obstructing the establishment of magnetic order and possibly giving the magnetic state liquid-like properties.
An important figure characterizing the degree of geometric frustration is the $f$-ratio [41]
$f=\theta_{W}/T_{c},$ (1)
where $\theta_{W}$ is the Weiss constant, the characteristic energy scale of interactions between the spins in the material, and $T_{c}$ is the critical temperature at which the spins order magnetically or undergo spin-glass freezing. On a non-frustrating or weakly frustrating lattice, the critical temperature $T_{c}$ is determined solely by the interaction strength in clean materials, and the $f$-ratio is of order unity. By contrast, strong geometric frustration prevents magnetic order from establishing and thus significantly lowers the ordering-transition temperature or possibly cancels the transition. Because the $f$-ratio quantifies a material’s ability to resist long-range order, it has been adopted as a figure-of-merit for the material’s potential to become a QSL.
Quenched disorder limits the $f$-ratio and, due to possibly inducing the spin-glass state, the material’s ability to become a QSL. In response to quenched disorder, GF magnets have recently been shown [55] to display several surprising, universal trends. Contrary to the common expectation that purifying a material suppresses spin-glass freezing, GF magnets exhibit the opposite trend: decreasing the density of vacancy defects, the most common type of disorder in such systems, increases the critical temperature of the spin-glass (SG) transition. For vanishingly low vacancy concentrations, a GF material would undergo a spin-glass transition at a “hidden energy scale” $T^{*}\lesssim\theta_{W}/10$, which remains finite despite the disappearance of the SG signal in susceptibility. While it remains difficult to fully eliminate quenched disorder, the purest systems exhibit robust signs of short-range order near $T^{*}$ in neutron scattering [7, 14, 53, 55] and specific heat, the quantity of interest for this work, regardless of the details of quenched disorder.
In this paper, we investigate the origin of the hidden energy scale and its relationship to the $f$-ratio. We demonstrate that the hidden energy scale is a property of the clean GF material and reveal its microscopic mechanism. We also argue that the $f$-ratio can be generalised to include the possibility that when a material undergoes a crossover at a low-temperature scale $T_{c}$, accompanied by thermodynamic deactivation of magnetic degrees of freedom, it does not exhibit a magnetic or spin-glass phase transition.
To illustrate the nature of the hidden energy scale, we give both experimental and theoretical evidence that the heat capacity $C(T)$ in GF materials exhibits two distinguishable peaks, with the lower-temperature peak located at temperatures of order $T^{*}$.
This lower-temperature peak comes from excitations similar to spin exchanges in chains of spins, whose energy is well separated from the energy of other excitations, similar to spin flips, which give rise to the higher-temperature peak in $C(T)$. We demonstrate that in the presence of quenched disorder, the material is sensitive to quenched disorder at temperatures $T\lesssim T^{*}$, which is consistent with the experimentally observed SG freezing of 3D GF materials at such temperatures. We also predict the entropy of the excitations giving rise to the lower-temperature peak, which allows for further verification of the proposed mechanism of the hidden energy scale.
Evidence for two distinct temperature scales in frustrated magnets. Key to the proposed scenario of the hidden energy scale is our observation, discussed in this paper, that GF magnets that lack long-range magnetic order at low temperatures universally exhibit two distinct temperature scales in the dependence of $C(T)$ on temperature.
In experiments [16, 44, 22, 34, 27, 6, 42, 2] on GF materials, in which the data extend over a large enough temperature range, these two temperature scales manifest themselves in two distinct, well-distinguishable peaks in $C(T)$. This behaviour is mirrored by numerical simulations [11, 60, 12, 33, 56, 50, 29, 30, 1, 54, 32, 48, 45, 38, 8, 31, 46, 19, 57] on small clusters of spins on frustrating lattices, exemplified by spins on the kagome lattice, whose heat capacity is shown in Fig. 1 (where the two peaks are labelled as $T^{*}$ and $\sim\theta_{W}$).
For some of the numerically simulated systems, the lower-temperature peak is less distinguishable from the larger, higher-temperature peak and forms a “shoulder” in its vicinity [54, 32, 8, 38, 45, 15]. We note that the resolution of numerical simulations may be limited by finite-size effects. Given, however, that the peaks are readily distinguishable in all experiments and given the preponderance of numerical evidence, the existence of two temperature scales in $C(T)$ is a universal feature of GF materials. In Table 1, we summarise the available experimental and numerical evidence [3].
Lattice System and references
Kagome $^{3}He$ on graphite [16, 22]
Numerics [11, 12, 1, 54, 32, 48, 45, 30, 50, 60, 29, 33, 56, 31, 57], this work
Triangular $NiGa_{2}S_{4}$ [34], $FeAl_{2}Se_{4}$ [27],
$NaYbO_{2}$ [6], $NaYbSe_{2}$ [42],
Numerics [8, 38, 46, 57]
Hyperkagome $Gd_{3}Ga_{5}O_{12}$ [44]
Numerics [19]
Spinel $CuGa_{2}O_{4}$ [2]
Table 1: Experimental and numerical evidence for two temperature scales in systems of spins on various geometrically frustrating lattices.
Figure 1: Heat capacity $C(T)$ of the spin-1/2 XXZXXZ model on a cluster of 18 sites on the kagome lattice computed from exact diagonalization with periodic boundary conditions for various values of the anisotropy parameter $\alpha$ [cf. Eq. (3a)]. The left inset shows $C(T)$ at very low temperatures, which may display a secondary low-temperature peak. The right inset shows the $\alpha$-dependence of the primary ($T^{*}$) and secondary ($T_{L}$) low-temperature peak positions. The existence of distinguishable low- and high-temperature peaks in $C(T)$ is in agreement with the results of Refs. [11, 12, 1, 54, 32, 48, 50, 60, 29, 33, 56, 31, 57] for the isotropic Heisenberg model ($\alpha=1$) and of Refs. [1, 57] for the XXZXXZ models. As we clarify in the Suppmental Material [4], the peaks may merge and turn into a plateau for $\alpha\sim 1$ and other boundary conditions.
Figure 2: Illustrations of Ising ground state hyridization via the transverse coupling $\hat{\cal H}_{\text{XY}}$ on the a) triangular, b) kagome, and c) pyrochlore lattices. Each picture shows two Ising ground states which are hybridized by $\hat{\cal H}_{\text{XY}}$. Fully-filled (open) circles indicate spin-up (down) on the corresponding sites, and are the same between the two states. Half-filled circles correspond to the upper spin orientations in one Ising state and the lower orientations in the other.
Excitations near the hidden energy scale. The higher energy scale in $C(T)$ is of order of the Weiss temperature
$\theta_{W}\sim ZJ$ (2)
in each material, where $Z$ and $J$ are, respectively, the coordination number of the magnetic sublattice of the material and the characteristic exchange coupling. The scale $\theta_{W}$ can be qualitatively understood as the energy of flipping a spin interacting with $Z$ neighbours with the interaction energy $J$. Although we focus, for simplicity, on spins-$1/2$ hereafter, our results hold for arbitrary spins and are consistent with the range of experimental examples.
To illustrate the origin of the lower energy scale, we consider the Hamiltonian
$\hat{\cal H}=\hat{\cal H}_{\text{Ising}}+\hat{\cal H}_{\text{XY}}+\hat{\cal H}_{\text{disorder}}+\hat{\cal H}_{\text{weak}}$ (3a)
$\hat{\cal H}_{\text{Ising}}=J\sum_{(ij)}{\hat{S}}_{i}^{z}{\hat{S}}_{j}^{z}$ (3b)
$\hat{\cal H}_{\text{XY}}=\alpha J\sum_{(ij)}\left({\hat{S}}_{i}^{x}{\hat{S}}_{j}^{x}+{\hat{S}}_{i}^{y}{\hat{S}}_{j}^{y}\right)$ (3c)
$\hat{\mathcal{H}}_{\mathrm{disorder}}=\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z}{\hat{S}}_{j}^{z}$ (3d)
of spins on a GF lattice, where the part $\hat{\cal H}_{\text{Ising}}+\hat{\cal H}_{\text{XY}}$ describes a Heisenberg $XXZ$ model; the summation is carried out over all pairs $(ij)$ of neighbouring spins; $J$ is the coupling strength between the spins’ $z$ components; $\hat{\cal H}_{\text{disorder}}$ is a perturbation created by weak quenched disorder whose exact form is not important but is considered, for concreteness, to come from the random fluctuations $\delta J_{ij}$ of the exchange couplings; $\hat{\cal H}_{\text{weak}}$ accounts for weak non-exchange, e.g. dipole-dipole, interactions with characteristic energies significantly smaller than $J$, and the anisotropy parameter $\alpha$ describes the relative strength of the transverse and longitudinal exchange interactions.
At $\alpha=0$ (the Ising limit) and $\hat{\cal H}_{\text{disorder}}=\hat{\cal H}_{\text{weak}}=0$, spins on a GF lattice have extensive degeneracy, i.e. degeneracy that scales exponentially with the system size, whereas single spin-flip excitations have a characteristic energy of the order of $\theta_{W}\sim JZ$.
For $0<\alpha\ll 1$, the degeneracy of the ground states is lifted by disorder and the transverse interactions, but the energies of respective low-lying excitations remain well separated from the characteristic energies of the excitations of the Ising model. Thus, at small $\alpha$ and weak disorder and other interactions, the model exhibits two types of excitations, adiabatically connected to, respectively, the ground states and the excited states of the Ising model, which gives rise to the two peaks in the heat capacity $C(T)$. We argue below that for most materials, the two peaks remain well separated for $\alpha\sim 1$, including in the isotropic Heisenberg models.
[…]
Examples of such “Ising instanton” processes are shown in Fig. 2. On the triangular and kagome lattices, $Z_{1}=1$, corresponding to the minimum-exchange processes between possible ground states showing in Figs. 2a and 2b. For the pyrochlore lattice, $Z_{1}=3$, corresponding to spin exchange in a loop of $6$ spins [17], as shown in Fig. 2c.
The characteristic eigenvalues of the operator $\hat{\tilde{H}}_{\text{Ising}}$ in Eq. (4) are given by $\theta_{W}\sim ZJ$, while the matrix elements of the operators $\hat{\cal H}_{\text{XY}}$ are given by $\alpha J/2$. The transition amplitude between the Ising ground states due to $Z_{1}$ pairwise spin-exchange processes can, therefore, be estimated as
$T^{*}\sim\left[\alpha J/\theta_{W}\right]^{Z_{1}-1}\alpha J\sim J\alpha^{Z_{1}}/Z^{Z_{1}-1}.$ (5)
The quantity (5) describes the characteristic energy splitting between the lowest energy levels of the $XXZ$ model with the Hamiltonian (3a)-(3c) in the limit of weak quenched disorder and weak non-exchange interactions. These states are adiabatically connected to the Ising ground states on the same lattice when the parameter $\alpha$ increases from 0 to a nonzero value. The quantity $Z_{1}$ in Eq. (5) is given by the minimum number of pairwise spin-exchange processes that can connect two Ising ground states. The value of $Z_{1}$, like $Z$, depends on the details of the lattice.
Excitations on the kagome lattice. Numerical simulations [1, 57] of the $XXZ$ Heisenberg models on the kagome lattice confirm the persistence of the two peaks up to $\alpha\sim 1$. To illustrate this trend, we show in Fig. 1 the heat heat capacity of a cluster of $18$ spins on the kagome lattice for various values of $\alpha$ that we obtain by exact diagonalization. The temperatures $T^{*}$ and $T_{L}$ of the two low-temperature peaks, shown in the inset in Fig. 1, scale $\propto\alpha$ at small $\alpha$, which demonstrates that the associated excitations at $\alpha\sim 1$ are adiabatically connected to single spin-exchange processes ($Z_{1}=1$) at $\alpha=0$ shown in Fig. 2b. We note the emergence of the small “secondary” peak at temperature $T_{L}$. Such a feature was also identified for the isotropic Heisenberg model ($\alpha=1$) in Refs. [50, 32, 48, 45, 31], where it was suggested to be a finite-size effect.
Constraint on the $f$-ratio. The scales (2) and (5) in a generic GF material determine the characteristic energies of the excitations that give rise to the two peaks in the behaviour of the heat capacity $C(T)$. As GF lattices typically have coordination numbers in the interval $Z=4\ldots 6$ and the quantity $Z_{1}=1\ldots 3$, the ratio
$f=\frac{\theta_{W}}{T^{*}}\sim\left(\frac{Z}{\alpha}\right)^{Z_{1}}$ (6)
of the two scales is large even when the parameter $\alpha$ is not small, resulting in a separation of the peaks of $C(T)$ in realistic materials. The order of magnitude of the ratio of the temperatures of higher and lower peaks, whose detailed analyses will be presented elsewhere, is consistent with the estimate (6).
At temperatures $T\lesssim T^{*}$, the excitations that give rise to the lower-temperature peak are suppressed, and the magnetic degrees of freedom of the material freeze. The hidden energy scale $T^{*}$, thus, constrains the $f$-ratio in GF magnets.
Spin-glass freezing in the presence of disorder. In 3D systems with quenched disorder, the thermodynamic freezing of the magnetic degrees of freedom at $T\lesssim T^{*}$ will lead to a spin-glass-freezing transition at temperatures $T_{c}\sim T^{*}$ for a broad range of disorder strengths.
To demonstrate this, we compute (see Supplemental Material [4] for details) the critical value of the variance $\varkappa_{c}=\left<\delta J_{ij}^{2}\right$> of the fluctuations $\delta J_{ij}$ of the exchange couplings in the mean-field replica-symmetric approximation [47, 5, 35] assuming that the fluctuations on different bonds are independent. Up to a non-universal prefactor that depends on the details of the lattice, the critical disorder strength of the glass transition is given by
$\varkappa_{c}(T)\sim\begin{cases}&\left[Z{\int}_{\begin{subarray}{c}\text{lower}\\ \text{peak}\end{subarray}}\frac{\rho(\xi)}{\xi^{2}}\,\mathrm{d}\xi\right]^{-1}=A(T^{*})^{2}/Z,\quad T\ll T^{*}\\ &T^{2}/Z,\quad T^{*}\ll T\ll\theta_{W},\end{cases}$ (7)
where the integral is taken over the low-energy states, i.e. states that give rise to the low-temperature peak of $C(T)$; $\rho(\xi)$ is the density of such states in the clean system, and $A$ is a dimensionless constant.
As we discuss in the Supplemental Material [4], the constant $A$ in Eq. (7) may be either zero or of order unity, depending on the excitation energy spectrum in a particular system, whose detailed investigation for specific materials we leave for future studies. In both cases, however, the critical disorder strength $\varkappa_{c}$ rapidly grows at temperatures $T\gtrsim T^{*}$. This suggests that the spin-glass transition will be observed at temperatures $T\sim T^{*}$ for a broad range of disorder strengths, in agreement with the recent analyses of spin-glass transitions in GF magnets [55].
The two-peak structure of heat capacity is specific to GF materials. We emphasise that non-GF materials will not display the two distinct energy scales we have described unless the spin-spin interactions contain two distinct energies [9]. Indeed, the ground states of a model that develops long-range magnetic order at low temperatures will not have extensive degeneracy. The excitations that are adiabatically connected to the Ising ground states when $\alpha$ is changed from $\alpha=0$ to a nonzero value will, therefore, have a vanishing contribution to $C(T)$ per spin and will not lead to the formation of the low-temperature peak.
Lattice Approximate entropy per site
Kagome $0.50183$ ([23, 52])
Triangular $0.323066$ ([59, 20, 24])
Pyrochlore $0.203$ ([39, 49, 25, 51, 28])
Table 2: Entropies of the ground states of the Ising model on various lattices.
Experimental test: entropy related to the “hidden energy scale”. In materials with well-separated peaks of $C(T)$ (see, e.g. Refs. [16, 34, 2]), the entropy $S_{\text{low}}=\int_{\begin{subarray}{c}\text{lower}\\ \text{peak}\end{subarray}}\frac{C(T)}{T}\mathrm{d}T$ associated with the lower peak is a well-defined quantity intrinsic to the pure material. Because the excitations that give rise to that peak are adiabatically connected to the ground states of the Ising model on the same lattice, the entropy $S_{\text{low}}$ matches the entropy of the ground states of that Ising model. The value of this entropy also remains unaltered by weak interactions and disorder so long as the peaks of $C(T)$ are well separated.
For common 22D lattices, the values of the respective Ising ground-state entropies can be obtained analytically, while for Ising models on other lattices, the entropy values are known from numerical simulations, as summarised in Table 2. Because the values of those (Ising) entropies match the entropies corresponding to the lower peaks in the more complicated, realistic model with the Hamiltonian (3a)-(3d), the value of that entropy of either peak of $C(T)$ can be used as an experimental test of the origin of the hidden energy scale.
This is exemplified by the layered triangular-lattice compound $NiGa_{2}S_{4}$, for which the lower peak entropy obtained from the digitized data of Ref. [34] is $S\approx 0.35$ per spin and is close to the Ising ground-state entropy $S_{\triangle}\approx 0.32$ for the triangular lattice [59, 20, 24]. If the low-temperature behaviour of $C(T)$ also exhibits a secondary low-temperature peak (cf. Fig. 1), the entropy under this secondary peak should be included in the respective entropy of the lower-temperature peak. The ratio $f^{\prime}\approx 8$ of the temperatures of the higher- and lower-temperature peaks in $NiGa_{2}S_{4}$ is close to the estimate $f=6$ given by Eq. (6).
Which peak is usually visible in experiments? Often, due to insufficient temperature range, only one of the two peaks is seen in experiments and it is not always clear which peak this is, the low- or the high-temperature one. Our thinking in this paper is focused on data for GF magnets containing $3d$ elements with $\theta_{W}$ values comparable to or greater than $100K$. As a result, there are few $C(T)$ measurements that capture this temperature region and the high-temperature peak (an exception is $NiGa_{2}S_{4}$, with $\theta_{W}=80K$ [34]).
By contrast, in materials with low $\theta_{W}$, such as rare-earth frustrated materials, the higher-temperature peak is usually observed. For example, in $YbMgGaO_{4}$ ($\theta_{W}$ = 4 K) [37] and $Ce_{2}Zr_{2}O_{7}$ ($\theta_{W}$ =0.6 K) [13], the peaks in $C(T)$ are observed, respectively, at $2K$ and $0.2K$. For such low temperatures of the higher-temperature peaks, the lower-temperature peaks should be expected at temperatures $T^{*}$ below the practical lower limit of most ultra-low-temperature experiments (around $50mK$).
Effect of the magnetic field. The low-temperature peak in the behaviour of $C(T)$ will be sensitive or insensitive to the external magnetic field, depending on the lattice and the range of interactions. For some lattices, such as the pyrochlore lattice [21], the “bilayer kagome” lattice of $SrCr_{8}Ga_{4}O_{19}$ [40] and the hyperkagome lattice of $Na_{4}Ir_{3}O_{8}$ [36, 18], the Ising ground states have zero magnetization if the exchange interactions between spins are nearest-neighbour. As a result, all the low-energy states of the Heisenberg models that are adiabatically connected to those Ising states have zero magnetizations. For this reason, the low-temperature peak of $C(T)$ in these materials is independent of the magnetic field, as observed in experiments [40, 36].
By contrast, if the Ising ground states allow for finite magnetizations, exemplified by the triangular and kagome lattices with nearest-neighbour interactions, the low-energy excitations of the corresponding Heisenberg models will, in general, have spins. Such excitations have been reported for the kagome lattice in Refs. [58, 50, 1, 26, 45, 57]. The low-temperature peaks in such materials will be sensitive to the magnetic field.
Conclusion. In conclusion, we argue, based on the analysis of available experimental and numerical data, that the temperature dependence of the heat capacity $C(T)$ exhibits two distinct temperature scales. The order of magnitude of the lower scale matches the “hidden energy scale” below which the cleanest of GF magnetic materials show spin-glass freezing and come from excitatinos adiabatically connected to the Ising ground states on the same lattice. The entropy associated with the lower-temperature peak matches the Ising ground-state entropy, which can be used to verify our predictions experimentally. As the value of the hidden energy scale is a property of the clean GF medium, it constrains the $f$-ratio, characterizing the degree of frustration.
Acknowledgments. We thank V. Elser for useful feedback on the manuscript and P. Prelovs̆ek and M. Ulaga for useful discussions. We thank an anonymous referee for bringing Ref. [27] to our attention. Our work has been supported by the NSF grant DMR2218130 and the Committee on Research at the University of California Santa Cruz.
Supplemental Material for
“Origin of the hidden energy scale and the $f$-ratio in geometrically frustrated magnets”
I Boundary conditions and the heat capacity of spins on the kagome lattice
In this section, we investigate the dependence of the numerical results for $C(T)$ on the boundary conditions of spin clusters. For the small clusters considered in Refs. [4, 5, 1, 20, 11, 18, 17, 9, 19, 23, 8, 12, 21, 10, 22] and in this work, the form of $C(T)$ is sensitive to the boundary conditions. This is exemplified by the two inequivalent $18$-site clusters of the kagome lattice shown in Fig. S1. The results we report in Fig. 1 were obtained using cluster (a), for which we obtain well-separated peaks in the heat capacity $C(T)$ in agreement with Refs. [1, 22]. However, while $XXZ$ Hamiltonians with small values of $\alpha$ on cluster (b) also display two well-separated peaks in $C(T)$, for larger values of $\alpha$, up through the isotropic Heisenberg case ($\alpha=1$), the high-temperature peak merges with the low-temperature one, forming a plateau rather than a distinct peak. These behaviours are shown in Fig. S2.
Figure S1: Two inequivalent $18$-site clusters on the kagome lattice. With periodic boundary conditions, cluster (a) has a threefold translation symmetry along the horizontal, while cluster (b) has a threefold translation symmetry along $60^{\circ}$ from the horizontal.
Figure S2: Heat capacity $C(T)$ (per spin) of the spin-$1/2$ XXZXXZ model on cluster (b) of Fig. S1. For large values of $\alpha$, there are not two well-separated peaks, in contrast to heat capacities on cluster (a) and in Refs. [1, 22].
While small spin clusters on the kagome lattice may not display good separation of the peaks in the behaviour of $C(T)$, the two peaks are rather distinct in the available experiments [6, 16, 7, 13, 3, 14, 2] on GF materials, which suggests that the separation improves in the limit of large system sizes, which are currently inaccessible by numerical methods. A more thorough understanding of the effects of boundary conditions and system size on the spectra and thermodynamic properties of kagome clusters awaits future study.
II Derivation of the critical disorder strength for glass freezing
In this section, we present a derivation of the phase boundary of the spin-glass state in the model of a GF magnet considered in the main text. For simplicity, we consider nearest-neighbour bond disorder in this model. The Hamiltonian of the system is given by
$\hat{\cal H}=\hat{\cal H}_{0}+\hat{\mathcal{H}}_{\mathrm{disorder}}$ (S1a)
$\hat{\mathcal{H}}_{\mathrm{disorder}}=\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z\gamma}{\hat{S}}_{j}^{z\gamma}$ (S1b)
where $\hat{\cal H}_{0}$ is the Hamiltonian of $R$ clean, independent replicas of the system; $\hat{\mathcal{H}}_{\mathrm{disorder}}$ describes quenched fluctuations of the $zz$-couplings and $\gamma$ is the replica index. In the main text, the disorder-free Hamiltonian of the system is given by $\hat{\cal H}_{\mathrm{Ising}}+\hat{\cal H}_{\mathrm{XY}}$ [see Eqs. (3b) and (3c)] in a single replica; however, the exact form of $\hat{\cal H}_{0}$ is not important for the present calculation. We consider each replica to consist of a finite lattice of $N$ sites with periodic boundary conditions and coordination number $Z$. The sum in Eq. (S1b) runs over all nearest-neighbor site pairs $(ij)$. Hereinafter, summation over repeated indices is implied.
Using the interaction representation, where we take Eq. (S1b) as a perturbation to $\hat{\cal H}_{0}$, we can write the partition function of the $R$ replicas for a particular realization of the bond fluctuations in the form
$Z^{R}[\{\delta J_{ij}\}]=Z_{0}^{R}\biggl{\langle}{\hat{T}}_{\tau}\,\mathrm{exp}\biggl{[}-\int_{0}^{\beta}\mathrm{d}\tau\,\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z\gamma}(\tau){\hat{S}}_{j}^{z\gamma}(\tau)\biggr{]}\biggr{\rangle}$ (S2)
where $\beta$ is the inverse temperature, $Z_{0}$ is the partition function of a single clean replica, ${\hat{S}}_{i}^{z\gamma}(\tau)$ is the spin operator in the interaction representation, ${\hat{T}}_{\tau}$ is the Matsubara-time-ordering operator, and the angle brackets denote equilibrium thermal averaging with respect to $\hat{\cal H}_{0}$.
Assuming, for simplicity, that the fluctuations are Gaussian-distributed independently on each bond with zero mean and variance $\varkappa$, we can average Eq. (S2) over disorder realizations to obtain
$\overline{Z}^{R}$ $=Z_{0}^{R}\biggl{\langle}{\hat{T}}_{\tau}\,\mathrm{exp}\biggl{[}\frac{\varkappa}{2}\sum_{(ij)}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,{\hat{S}}_{i}^{z\gamma}(\tau){\hat{S}}_{j}^{z\gamma}(\tau){\hat{S}}_{i}^{z\delta}(\tau^{\prime}){\hat{S}}_{j}^{z\delta}(\tau^{\prime})\biggr{]}\biggr{\rangle}$
$=Z_{0}^{R}\biggl{\langle}{\hat{T}}_{\tau}\,\mathrm{exp}\biggl{[}\frac{\varkappa}{4}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,{\hat{S}}_{i}^{z\gamma}(\tau){\hat{S}}_{i}^{z\delta}(\tau^{\prime})K_{ij}{\hat{S}}_{j}^{z\gamma}(\tau){\hat{S}}_{j}^{z\delta}(\tau^{\prime})\biggr{]}\biggr{\rangle},$ (S3)
where in writing the second line of Eq. (S3) we have introduced the matrix K{K} whose entries are $K_{ij}=1$ if sites $i$ and $j$ are first neighbors, and 0 otherwise.
We now decouple the quartic product of spin operators in the exponent of Eq. (S3) using the Hubbard-Stratonovich transformation:
$\overline{Z}^{R}=Z_{0}^{R}\int{\cal D}q\,\mathrm{exp}\biggl{[}-\frac{\varkappa}{4}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,q_{i\tau\tau^{\prime}}^{\gamma\delta}K_{ij}q_{j\tau\tau^{\prime}}^{\gamma\delta}\biggr{]}\biggl{\langle}{\hat{T}}_{\tau}\,\mathrm{exp}\biggl{[}\frac{\varkappa}{2}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,q_{i\tau\tau^{\prime}}^{\gamma\delta}K_{ij}{\hat{S}}_{j}^{z\gamma}(\tau){\hat{S}}_{j}^{z\delta}(\tau^{\prime})\biggr{]}\biggr{\rangle}.$ (S4)
To evaluate the thermal average in Eq. (S4), we assume that the glass transition is of the second order and consider the system close to the transition. Hence, the glass order parameter is small. Employing a replica-symmetric, mean-field ansatz for the order parameter $q_{i\tau\tau^{\prime}}^{\gamma\delta}\rightarrow q$, we expand the term in the angular brackets in Eq. (S4) as
$\quad\biggl{\langle}{\hat{T}}_{\tau}\,\mathrm{exp}\biggl{[}\frac{\varkappa q}{2}\sum_{(\gamma\delta)}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,K_{ij}{\hat{S}}_{j}^{z\gamma}(\tau){\hat{S}}_{j}^{z\delta}(\tau^{\prime})\biggr{]}\biggr{\rangle}$
$=1+\biggl{\langle}{\hat{T}}_{\tau}\,\frac{\varkappa q}{2}\sum_{i,(\gamma\delta)}K_{ij}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,{\hat{S}}_{j}^{z\gamma}(\tau){\hat{S}}_{j}^{z\delta}(\tau^{\prime})\biggr{\rangle}$
$+\biggl{\langle}{\hat{T}}_{\tau}\,\frac{\varkappa^{2}q^{2}}{4}\sum_{ik,(\gamma_{1}\delta_{1}),(\gamma_{2}\delta_{2})}K_{ij}K_{k\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,\mathrm{d}\theta\,\mathrm{d}\theta^{\prime}\,{\hat{S}}_{j}^{z\gamma_{1}}(\tau){\hat{S}}_{j}^{z\delta_{1}}(\tau^{\prime}){\hat{S}}_{\ell}^{z\gamma_{2}}(\theta){\hat{S}}_{\ell}^{z\delta_{2}}(\theta^{\prime})\biggr{\rangle}+{\cal O}(q^{3}),$ (S5)
where the now explicit replica sums run over all pairs $(\gamma\delta)$ of replicas. The ${\cal O}(q)$ term in Eq. (S5) vanishes since clean replicas are independent of one another and the Hamiltonian $\hat{\cal H}_{0}$ is assumed to have global spin inversion symmetry. For the ${\cal O}(q^{2})$ term, the replica pair sums give two types of contributions: those in which either $\gamma_{2}=\gamma_{1}$, $\delta_{2}=\delta_{1}$ or $\gamma_{2}=\delta_{1}$, $\delta_{2}=\gamma_{1}$ and those in which there are three or four distinct replica indices. The latter type of term vanishes for the same reason as the ${\cal O}(q)$ contribution. The former type of term, however, does not vanish, and yields
$\frac{1}{4}R(R-1)\varkappa^{2}\beta^{2}q^{2}Z^{2}N\sum_{\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,g(0,0;\ell,\tau)g(0,0;\ell,\tau^{\prime}),$ (S6)
where we define
$g(j,\tau;\ell,\theta)\equiv\bigl{\langle}{\hat{T}}_{\tau}\,{\hat{S}}_{j}^{z\gamma}(\tau){\hat{S}}_{\ell}^{z\gamma}(\theta)\bigr{\rangle}.$ (S7)
With this ansatz, we then have, to the order ${\cal O}(q^{2})$,
$\overline{Z}^{R}=Z_{0}^{R}+\frac{Z_{0}^{R}}{4}R(R-1)\varkappa\beta^{2}q^{2}ZN\biggl{[}\varkappa Z\sum_{\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,g(0,0;\ell,\tau)g(0,0;\ell,\tau^{\prime})-\frac{1}{2}\biggr{]}.$ (S8)
We then obtain the disorder-averaged free energy functional for the glass transition to ${\cal O}(q^{2})$ as
$\beta{\cal F}(q)=\underset{R\rightarrow 0}{\mathrm{lim}}\frac{\overline{Z}^{R}-1}{R}=\frac{1}{4}\varkappa\beta^{2}q^{2}ZN\biggl{[}\frac{1}{2}-\varkappa Z\sum_{\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,g(0,0;\ell,\tau)g(0,0;\ell,\tau^{\prime})\biggr{]}+{\cal O}(q^{3})$ (S9)
which yields the critical disorder strength for glass freezing:
$\varkappa_{c}^{-1}(\beta)=2Z\sum_{\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,g(0,0;\ell,\tau)g(0,0;\ell,\tau^{\prime}).$ (S10)
II.1 Critical disorder strength in the effective Ising regime
In what immediately follows, we analyse the phase boundary (S10) in the temperature interval $T^{*}\ll T\ll\theta_{W}$. For simplicity, to capture the behaviour of the magnetic materials qualitatively correctly, it is sufficient to consider the case of small $\alpha\ll 1$.
For describing the contribution of the excitations with energies $T^{*}\ll E\ll\theta_{W}$ in Eq. (S10), we can neglect the small transverse spin-exchange-induced splitting between states adiabatically connected to Ising ground states and consider the respective states to be approximately equally thermally populated. Quantum fluctuations are thus washed out, and the system is in the effective Ising regime. Also, the excitations of the Ising Hamiltonian with energies of order $\theta_{W}$ can be neglected.
We can write the $\tau$-integral over the correlation function (S7) in the form
$\int_{0}^{\beta}\mathrm{d}\tau\,g(0,0;\ell,\tau)=\frac{1}{Z_{0}(\beta)}\sum_{\bf k}\sum_{{\bf k}^{\prime}}\frac{\mathrm{e}^{-\beta\xi_{{\bf k}}^{\prime}}-\mathrm{e}^{-\beta\xi_{\bf k}}}{\xi_{\bf k}-\xi_{{\bf k}^{\prime}}}\langle{\bf k}|{\hat{S}}_{\ell}^{z}|{\bf k}^{\prime}\rangle\langle{\bf k}^{\prime}|{\hat{S}}_{0}^{z}|{\bf k}\rangle,$ (S11)
where $\{|{\bf k}\rangle\}$ and $\{\xi_{\bf k}\}$ are respectively the eigenstates and eigenenergies of a single clean replica (we have dropped the replica indices, as they are no longer important). As discussed above, both the clean partition function $Z_{0}(\beta)$ and the sum in Eq. (S11) will be dominated by the low-energy excitations adiabatically connected to the ground state manifold of the corresponding Ising model. In particular, the partition function can be approximated as $Z_{0}(\beta)\approx d_{\mathrm{G}}$, where $d_{G}$ is the ground state degeneracy of the Ising system, and we can make the approximation
$\frac{\mathrm{e}^{-\beta\xi_{{\bf k}}^{\prime}}-\mathrm{e}^{-\beta\xi_{\bf k}}}{\xi_{\bf k}-\xi_{{\bf k}^{\prime}}}\approx\beta.$ (S12)
Hence, Eq. (S10) reduces to
$\varkappa_{c}^{-1}(\beta)\approx\frac{2Z\beta^{2}}{d_{\mathrm{G}}^{2}}\sum_{\ell}\biggl{(}\sum_{{\bf k}_{0}}\sum_{{\bf k}_{0}^{\prime}}\langle{\bf k}_{0}|{\hat{S}}_{\ell}^{z}|{\bf k}_{0}^{\prime}\rangle\langle{\bf k}_{0}^{\prime}|{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle\biggr{)}^{2},$ (S13)
where the subscript 0 on the state indices denotes a state which is adiabatically connected to an Ising ground state. We can decompose Eq. (S13) further as follows: writing
$\sum_{{\bf k}_{0}^{\prime}}|{\bf k}_{0}^{\prime}\rangle\langle{\bf k}_{0}^{\prime}|=\hat{\mathbb{I}}-\hat{\mathcal{P}}_{E}$ (S14)
where $\hat{\mathcal{P}}_{E}$ projects onto the excitations of $\hat{\cal H}_{0}$ which are adiabatically connected to the excited states of the Ising model, Eq. (S13) becomes
$\varkappa_{c}^{-1}(\beta)$ $\approx\frac{2Z\beta^{2}}{d_{G}^{2}}\biggl{(}d_{G}^{2}-2d_{G}\sum_{{\bf k}_{0}}\langle{\bf k}_{0}|{\hat{S}}_{0}^{z}\hat{\mathcal{P}}_{E}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle+\biggl{[}\sum_{{\bf k}_{0}}\langle{\bf k}_{0}|{\hat{S}}_{0}^{z}\hat{\mathcal{P}}_{E}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle\biggr{]}^{2}$
$+\sum_{\ell>0}\biggl{[}\biggl{\{}\sum_{{\bf k}_{0}}\langle{\bf k}_{0}|{\hat{S}}_{\ell}^{z}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle\biggr{\}}^{2}-2\sum_{{\bf k}_{0}}\sum_{{\bf k}_{0}^{\prime}}\langle{\bf k}_{0}|{\hat{S}}_{\ell}^{z}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle\langle{\bf k}_{0}^{\prime}|{\hat{S}}_{0}^{z}\hat{\mathcal{P}}_{E}{\hat{S}}_{0}^{z}|{\bf k}_{0}^{\prime}\rangle+\biggl{\{}\sum_{{\bf k}_{0}}\langle{\bf k}_{0}|{\hat{S}}_{\ell}^{z}\hat{\mathcal{P}}_{E}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle\biggr{\}}^{2}\biggr{]}\biggr{)}$
$=2Z\beta^{2}\biggl{(}1+\frac{1}{d_{G}^{2}}\sum_{\ell>0}\biggl{[}\sum_{{\bf k}_{0}}\langle{\bf k}_{0}|{\hat{S}}_{\ell}^{z}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle\biggr{]}^{2}\biggr{)},$ (S15)
where in writing the last line of Eq. (S15) we have used that since Ising ground states are eigenstates of ${\hat{S}}_{0}^{z}$, and $\hat{\mathcal{P}}_{E}$ projects onto superpositions of Ising excited states, $\hat{\mathcal{P}}_{E}{\hat{S}}_{0}^{z}|{\bf k}_{0}\rangle$ vanishes. The second term in the last line of Eq. (S15) is a lattice-dependent factor of order unity. Thus, in the effective Ising regime, $\varkappa_{c}^{-1}\sim 2Z/T^{2}$.
II.2 Critical disorder strength at very low temperatures
We now consider Eq. (S10) at low temperatures, $T\ll T^{*}$. In this limit, Eq. (S11) reduces to
$2\sum_{{\bf k}}\frac{1}{\xi_{{\bf k}}}\mathrm{Re}[\langle{\bf k}|{\hat{S}}_{\ell}^{z}|0\rangle\langle 0|{\hat{S}}_{0}^{z}|{\bf k}\rangle],$ (S16)
assuming the Hamiltonian $\hat{\cal H}_{0}$ has a non-degenerate ground state.
Focusing on the matrix element $\langle{\bf k}|{\hat{S}}_{\ell}^{z}|0\rangle$, it can be shown that the many-body states of an $N$-site lattice with periodic boundary conditions take the form
$\mathrm{exp}\biggl{(}\frac{i}{N}{\bf k}\cdot\sum_{j=1}^{N}{\bf R}_{j}\biggr{)}|u\rangle,$ (S17)
similarly to Bloch’s states in the single-particle case, where $|u\rangle$ is a state invariant under simultaneous translation of all lattice sites by a lattice vector. Hence, defining $^{\hat{T}}_{\ell}$ as the translation operator which translates states by the lattice vector $\bf R$$_{\ell}$ connecting sites $\ell$ and 0, we have
$\langle{\bf k}|{\hat{S}}_{\ell}^{z}|0\rangle=\langle{\bf k}|{\hat{T}}_{\ell}{\hat{S}}_{0}^{z}{\hat{T}}_{\ell}^{\dagger}|0\rangle=\mathrm{e}^{i{\bf k}\cdot{\bf R}_{\ell}}\langle{\bf k}|{\hat{S}}_{0}^{z}|0\rangle.$ (S18)
Hence, for Eq. (S10), we have
$\varkappa_{c}^{-1}=8Z\sum_{\ell}\sum_{{\bf k}}\sum_{{\bf k}^{\prime}}\frac{\mathrm{e}^{i{\bf R}_{\ell}\cdot({\bf k}-{\bf k}^{\prime})}}{\xi_{{\bf k}}\xi_{{\bf k}^{\prime}}}|\langle{\bf k}|{\hat{S}}_{0}^{z}|0\rangle|^{2}|\langle{\bf k}^{\prime}|{\hat{S}}_{0}^{z}|0\rangle|^{2}=8ZN\sum_{{\bf k}}\frac{1}{\xi_{{\bf k}}^{2}}|\langle{\bf k}|{\hat{S}}_{0}^{z}|0\rangle|^{4}.$ (S19)
Now, since the non-vanishing contributions to the expression (S19) are from the states $\{|{\bf k}\rangle\}$ that are superpositions of Ising ground states, whose degeneracy scales exponentially with $N$ in GF systems, the quantity $|\langle{\bf k}|{\hat{S}}_{0}^{z}|0\rangle|^{2}$ admits an expansion in powers of $N^{-1}$ with coefficients of order unity for each $\bf k$}. Thus, in the thermodynamic limit, the critical disorder strength is given by
$\varkappa_{c}^{-1}\sim\frac{8Z}{N}\sum_{{\bf k}}\frac{1}{\xi_{{\bf k}}^{2}}\rightarrow 8Z\int\frac{\rho(\xi)}{\xi^{2}}\,\mathrm{d}\xi$ (S20)
where again the integral in Eq. (S20) is taken over the excitations which are connected to the Ising ground state manifold, i.e. the states giving rise to the low-temperature peak in the specific heat.
The critical disorder strength $\varkappa_{c}$ may be zero or finite depending on whether the integral in the right-hand-side of Eq. (S20) diverges or converges. The convergence of this integral is determined by the scaling of the low-energy density of states in a particular system and by whether the excitations are gapped in the thermodynamic limit.
If the excitations remain gapped in the thermodynamic limit, the integral converges, and $\varkappa_{c}$ remains finite as $T\rightarrow 0$. If, however, the excitations are gapless, the integral will converge for $\rho(\xi)$ scaling faster than linearly at small $\xi$, and diverge otherwise. This means that in a gapless 33D system with the power-law excitation dispersion $\xi(k)\propto k^{p}$ in the long-wave limit $k\rightarrow 0$,
$\varkappa_{c}(T=0)=\begin{cases}&0,\quad p\geq 3/2\\ &A(T^{*})^{2}/Z,\quad p<3/2\end{cases}$ (S21)
where $A$ in the second line is a lattice-dependent factor of order unity.
Equation (S21) suggests that in 3D systems with linearly dispersive excitations, which are expected, e.g., on the pyrochlore lattice [15], and in systems with a constant density of states, the critical disorder strength $\varkappa_{c}(T=0)$ at zero temperature is finite. By contrast, quadratic dispersion will lead to the vanishing of $\varkappa_{c}(T=0)$. Detailed analyses of the critical disorder strength for particular materials will be presented in future work.